A company making tires for bikes is concerned about the exact width of their cyclocross tires. The company has a lower specification limit of 23 mm and an upper specification limit of 23.6 mm. The standard deviation is 0.2 mm and the mean is 23.3 mm Round your answer to 4 decimal places.) What is the process capability index for the process?
Answer #1

Upper specification limit (USL) = 23.6 mm

Lower specification limit (LSL) = 23mm

Standard deviation (Sigma) = 0.2 mm

mean = 23.3 mm

Process capability index = (USL - LSL) / 6Sigma

= (23.6 - 23.0) / (6 x 0.2)

= 0.6/1.2 = 0.5